undeadknight26 (undeadknight26):

How do i turn a inequality into a point slope form equation and solve it?

OpenStudy (solomonzelman):

y=mx+b put your equation into that form and plug in values for x to graph it.

OpenStudy (solomonzelman):

Do you have a specific question?

undeadknight26 (undeadknight26):

y ≥ (-2/3)x -3

undeadknight26 (undeadknight26):

so is it : -2/3x -3 = y?

OpenStudy (solomonzelman):

No

OpenStudy (solomonzelman):

But you can use -2/3x -3 = y? to find the points.

OpenStudy (solomonzelman):

I am plugging in -3 and -6 for x, if x=-3 y= -1 so the first coordinate would be (3, -1) if x=-6 y= 1 so the second coordinate would be (6, 1) graph it.

OpenStudy (solomonzelman):



undeadknight26 (undeadknight26):

where did u get 6 and 1?

OpenStudy (solomonzelman):

and shade the area that is less than y or equal to it (incl to the line)  and the points are if x=-3 y= -1 (3, -1) if x=-6 y= 1 (6, 1) I used y = (-2/3)x -3 to get the points and then graphed it as y ≥ (-2/3)x -3 y ≥ meaning that the graph is everything that is less than or equal to that line. Still lost?